Could someone please help me with a few problems on cramer's rule? I want to know if I'm doing it right.Thank you so much in advance.
1)2x + 5y = 3
x + 4y = 2
2)x + 2y = 4
3x - 2y = -12

Answer by nerdybill(5413)

This site explains Cramer's rule in detail:
http://www.purplemath.com/modules/cramers.htm
1)2x + 5y = 3
x + 4y = 2
Determinant of D:
2 5
1 4
equals (2*4)-(1*5) = 8-5 = 3
Determinant of D(x):
3 5
2 4
equals (3*4)-(2*5) = 12-10 = 2
x = D(x)/D = 2/3
Determinant of D(y):
2 3
1 2
equals (2*2)-(1*3) = 4-3 = 1
y = D(y)/D = 1/3


2)x + 2y = 4
3x - 2y = -12
Determinant of D:
1 2
3 -2
equals (1*(-2))-(3*2) = -2-6 = -8
Determinant of D(x):
4 2
-12 -2
equals (4*(-2))-((-12)*2) = -8-(-24) = 16
x = D(x)/D = 16/(-8) = -2
Determinant of D(y):
1 4
3 -12
equals (1*(-12))-(4*3) = -12-12 = -24
y = D(y)/D = -24/(-8) = 3